the barangay captain wants to conduct a study to estimate average playing time of children ages 7 to 12 he wants to be 99% confident that the sample mean will be within 15 minutes of the true mean. If he can approximate a population standard deviation by 40 minutes and assume an approximately normal distribution, how large a sample should he get?

1
Expert's answer
2022-05-12T14:45:56-0400

The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} = 2.5758."


"E=z_c\dfrac{\sigma}{\sqrt{n}}=15/2"

"n\ge(\dfrac{2.5758(40)}{7.5})^2"

"n=189"